Jolly Juice Ltd. is expected to pay a $3.00 dividend next year and a $4 dividend in two years. After that, dividends are expected to grow at an annual rate of 5% forever. If investors require a return of 10% on the investment, what should jolly juice shares sell for today?
Answer #1

Step 1: Computation of market price at the end of year 2 using Gordon Growth Model

Using Gordon Growth Model

P2 = D3 / (Ke – g)

Where,

P2 - Market price at the end of year 2 = ?

D3 - Expected dividend in year 3 = D2*(1+g) = 4*1.05 = 4.20

Ke – Cost of equity = 10%

G – Growth rate in dividend = 5%

P2 = 4.2/(.10-.05)

= 4.2/.05

= $84.00

Step 2: Computing current share price by discounting the cashflow at required return

Year Dividend PVF@10% Present Value (Cashflow*PVF)
1 3                   0.909 2.73
2 88 (4+84)                   0.826 72.73

Current share price = \sumCashflow*PVF

= 2.73+72.73

= $75.45

You can use the equation 1/(1+i)^n to find PVF using calculator